please suggest me a way to solve maxima minima problems faster without differentiating which makes it long...and time consuming..........

Grade:12

2 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear sohan

You can't think maxima minima problem without differentiation

in most of the problem you need to use differentiation to find mox or mi vaue

in few cases u can solve problem by using direct result.

for exm.

 if you want to know max/min value of 2 degree polynomial then u can find it by knowing the vertex of parabola.

 why u r  worring about differentiation ,it is oe of the simplest topic in maths.

 


Please feel free to post as many doubts on our discussion forum as you can. If you find any question Difficult to understand - post it here and we will get you the answer and detailed solution very quickly.
 We are all IITians and here to help you in your IIT JEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin

 

Askiitians_Expert Yagyadutt
13 years ago

Hii soham

 

Well it depends on question dear...whether the question can be reduced in some format where differntiation is required or not...!

 

Like ...f(x) = x^2 - 2x + 3    Mind the minimum value 

 

f(x) = x^2 - 2x + 1 + 2  = (x-1)^2 + 2 

 

(x-1)^2 is always positive with minimum value zero....so f(x) has minimum value = 2

 

In some question like ....F(x) = x + 1/x  where x>0 ..find the minimum value...then you need to apply AM>=GM ..it will be easier..

 

So there are some class of question which can be done without using Differentiation...but only if you are able to determine it...Other wise ..differentiation is the only way to solve for maximum and minimum value of function..

Like in these question

 

F(x) = 3x^3 + 5 x^2 + 3x + 6  you have to differntiate it to find maximum and minimum value..

 

Sometime in trignometry Question....differentiation is not required

Like f(x) = 3sinx + 4cosx ...

....But in general you must know the way of finding maximum and minimum using differentiation

 

Best of luck

Yagya

askiitians_expert